Jan is 7 years younger than Lucy and 4 years older than Sam The average of their ages is 16. How old is Jan? Sam? Lucy?
Answer by lwsshak3(11628)

Jan is 7 years younger than Lucy and 4 years older than Sam The average of their ages is 16. How old is Jan? Sam? Lucy?
***
let x=Jan's age
x+7=Lucy's age
x-4=Sam's age
..
x+(x+7)+(x-4)/3=16
3x+3=48
3x=45
x=15
x+7=22
x-4=11
Jan's age=15
Lucy's age=22
Sam's age=11